# The reference interval in higher-order stochastic dominance

### v3

@@ 1 Introduction
Stochastic dominance is a widely used concept in economics, finance, and decision-making under uncertainty, providing a robust method for comparing distributions of uncertain outcomes. This concept is essential in evaluating risk preferences without relying on a specific utility function or preference model, which allows for broad applications across various fields (Levy (2015); Shaked and Shanthikumar (2007)).
@@
Despite its widespread use, the definition of higher-order stochastic dominance lacks consistency across the literature, sometimes leading to interpretational ambiguity. Consider, for example, two distributions, $F$ and $G$, each supported over the interval $[0,1]$. At first glance, one might assume that the question of whether $F$ dominates $G$ in fourth-order stochastic dominance would yield a straightforward answer. However, the consequence can depend significantly on the choice of the evaluation interval. For instance, if we assess the dominance using only the interval $[0,1]$, $F$ may not dominate $G$. Yet, extending the interval to $[0,2]$ might flip the assessment, resulting in $F$ dominating $G$ [^1] For a detailed discussion, see Example 1, where the specific distributions of $F=(8/9+\epsilon)\delta_{2/9}+(1/9-\epsilon)\delta_{1}$ and $G=\delta_{0}/3+2\delta_{4/9}/3$ with $\epsilon=1/100$. This highlights a crucial aspect of higher-order stochastic dominance: It can vary with alterations in the interval considered. This issue has led to ambiguous formulations of higher-order stochastic dominance across various texts. For example, Definition 7 in Baiardi et al. (2020) and the related definitions in Section 2.3 of Denuit and Eeckhoudt (2010) both adopted an arbitrary interval that encompasses the support of the distribution, but the definition actually depends on the choice of the interval.
@@
We say that the two formulations are consistency if the ranking between distributions $F$ and $G$ supported in $[a,b]$ remains the same when assessed under $n{\rm SD}_{\mathbb{R}}$ or $n{\rm SD}_{[a,b]}$. To the best of our knowledge, although various papers hint at the inconsistency issue under different settings (see Section 4 for a review), the consistency of the two formulations of higher-order stochastic dominance was explicitly discussed only in Fang and Post (2022). In their Section 2.2, they contended that $n{\rm SD}_{[a,b]}$ imposes a more stringent criterion than $n{\rm SD}_{\mathbb{R}}$, suggesting that inconsistencies might arise when $n\geq 4$. We formally encapsulate these observations in our Proposition 1, providing a detailed analysis and illustrating the inconsistencies for cases where $n\geq 4$ with a straightforward counterexample in Example 1.
@@
One implication of our results is that, since stochastic domination relations (with $n\geq 4$) get strictly stronger when the reference interval shrinks, it affects both their applications and characterization results. For instance, a stochastic dominance relation is easier to hold when we enlarge the reference interval, which are usually harmless for real-data applications. The results also illustrate a drawback of the higher-order stochastic dominance relations: As stochastic dominance is mostly used as a robust tool for ordering risks without assuming specific preferences, the fact that they depend on a reference interval — a subjective choice of the modeler — jeopardizes their robustness interpretation.
The rest of the paper is organized as follows. Section 2 introduces notation and contains all main results. All proofs are presented in Section 3. In Section 4, we present a literature review on stochastic dominance when applied to distributions restricted to specific subsets of $\mathbb{R}$.
@@ 2 Main results
For $a,b\in[-\infty,\infty]$ with $a<b$, denote by $\mathcal{X}_{[a,b]}$ the set of all bounded random variables taking values in $[a,b]$. For simplicity, we write $\mathcal{X}:=\mathcal{X}_{[-\infty,\infty]}$. We use capital letters, such as $X$ and $Y$, to represent random variables, and $F$ and $G$ for distribution functions. For $X\in\mathcal{X}$, we write $\mathbb{E}[X]$ for the expectation of $X$. Let $F_{X}$ denote the distribution function of $X$. We use $\delta_{\eta}$ to represent the point-mass at $\eta\in\mathbb{R}$. For a real-valued function $f$, let $f_{-}^{\prime}$ and $f^{\prime}_{+}$ be the left and right derivative of $f$, respectively, and denote by $f^{(n)}$ the $n$th derivative for $n\in\mathbb{N}$. Whenever we use the notation $f_{-}^{\prime}$, $f^{\prime}_{+}$ and $f^{(n)}$, it is understood that they exist. Denote by $[n]:=\{1,\dots,n\}$ with $n\in\mathbb{N}$. In this paper, all terms like “increasing”, “decreasing”, “convex”, and “concave” are in the weak sense.
@@
It is well-known that $F_{X}^{[n]}(\eta)$ is connected to the expectation of $(\eta-X)_{+}^{n}$ (see e.g., Proposition 1 of Ogryczak and Ruszczyński (2001)):
@@
[Shaked and Shanthikumar (2007)] Let $n\in\mathbb{N}$. For $X,Y\in\mathcal{X}$, we say that $X$ dominates $Y$ in the sense of $n$th-order stochastic dominance on $\mathbb{R}$ ($n{\rm SD}_{\mathbb{R}}$), denoted by $X\geq_{n}Y$ or $F_{X}\geq_{n}F_{Y}$ if
@@
[Jean (1980)] Let $a,b\in[-\infty,\infty]$ with $a<b$ and $n\in\mathbb{N}$. For $X,Y\in\mathcal{X}_{[a,b]}$, we say that $X$ dominates $Y$ in the sense of $n$th-order stochastic dominance on $[a,b]$ ($n{\rm SD}_{[a,b]}$), denoted by $X\geq_{n}^{[a,b]}Y$ or $F_{X}\geq_{n}^{[a,b]}F_{Y}$ if
@@
$\mathbb{E}[(\eta-X)_{+}^{n-1}]\leq\mathbb{E}[(\eta-Y)_{+}^{n-1}],~{}\forall\eta\in[a,b]~{}{\rm and}~{}\mathbb{E}[(b-X)^{k}]\leq\mathbb{E}[(b-Y)^{k}]~{}{\rm for}~{}k\in[n-1].$
For $n\in[4]$, $n$SD corresponds to the well-known first, second, third and fourth-order stochastic dominance. Risk aversion, which includes aversion to mean-preserving spreads, aligns with second-order stochastic dominance as described in Rothschild and Stiglitz (1970). Higher orders of stochastic dominance, specifically third and fourth orders, cater to decision makers with more refined risk preferences. Third-order stochastic dominance reflects prudence (Kimball (1990)), while fourth-order dominance corresponds to temperance (Kimball (1992)). The characterizations of these preferences, as detailed by Eeckhoudt and Schlesinger (2006), extend from the traditional concept of mean-preserving spreads to a broader framework of risk apportionment.
For two random variables $X,Y\in\mathcal{X}_{[a,b]}$, both $n{\rm SD}_{\mathbb{R}}$ and $n{\rm SD}_{[a,b]}$ can be used to rank their order. This raises a natural question of whether these ranking relations are consistent. To examine this, we define two sets of utility functions that are regular $n$-increasing concave (Denuit and Eeckhoudt (2013)) over different domains:222 From the definition in Liu (2014), the utility functions in $\mathcal{U}_{n}$ exhibit $k$th-degree risk aversion for $k\in[n]$. Letting $n\to\infty$, $\mathcal{U}_{\infty}$ becomes the set of all completely monotone functions, which is well studied in the mathematics literature and is closely linked to Laplace–Stieltjes transforms (see, e.g., Schoenberg (1938)). In this case, utility functions express mixed risk aversion, as discussed in Caballé and Pomansky (1996).
@@
Below, we present a specific example involving two random variables $X,Y\in\mathcal{X}_{[0,1]}$. This example demonstrates that $X\geq_{4}Y$ and $X\not\geq_{4}^{[0,1]}Y$, serving as a counterexample to the implication (iii) $\Rightarrow$ (i) in Proposition 1.
@@
It holds that
@@
Hence. $X\not\geq_{4}^{[0,1]}Y$.
@@
By standard calculation, one can check that $\mathbb{E}[\left(\eta-X\right)_{+}^{3}]\leq\mathbb{E}[\left(\eta-Y\right)_{+}^{3}]$ for all $\eta\in\mathbb{R}$ (see Figure 1 for an intuitive illustration), which implies that $X\geq_{4}Y$. Moreover,
$\mathbb{E}\left[\left(2-X\right)^{2}\right]-\mathbb{E}\left[\left(2-Y\right)^{2}\right]=-\frac{13}{2916}<0.$
Hence, we also have $X\geq_{4}^{[0,2]}Y$.
@@ 3.1 Proof of Proposition 1
Note that $\overline{\mathcal{U}}_{n}^{[a,b]}$ is the closure of ${\mathcal{U}}_{n}^{[a,b]}$ with respect to pointwise convergence. We only need to consider the statements (i), (ii), (iv) and (v). Below we will prove Proposition 1 by three steps.
@@
In Step (b), the equivalence of (i) and (ii) was well-established, as discussed in Section 2. Specifically, the implication (i) $\Rightarrow$ (ii) can be shown by using integration by parts, greatly facilitated by the condition $F_{X}^{[k]}(b)\leq F_{Y}^{[k]}(b)$ for $k\in[n]$ derived from $X\geq_{n}^{[a,b]}Y$. This condition, however, is absent in the analysis of the implication (iv) $\Rightarrow$ (v). In fact, the implication (iv) $\Rightarrow$ (v) can be verified by the insight that every $u\in\mathcal{U}_{n}$ is a positive linear combination of singularity functions in the set $\{f:x\mapsto-(\eta-x)_{+}^{n-1}|\eta\in\mathbb{R}\}$; see Williamson (1956). We provide a self-contained proof based on the integration by parts.
Before showing the proofs, we present an auxiliary lemma for later use.
@@
(iv) $\Rightarrow$ (i): It follows from Lemma 1 that (iv) implies $\mathbb{E}[X]\geq\mathbb{E}[Y]$. This completes the proof. ∎
@@
The implication (i) $\Leftrightarrow$ (ii) has been verified in Theorem 1 of Eeckhoudt et al. (2009). The implication (ii) $\Rightarrow$ (v) is trivial. The implication (v) $\Rightarrow$ (iv) is supported by the fact that the mapping $x\mapsto-(\eta-x)_{+}^{n-1}$ is contained in $\overline{\mathcal{U}}_{n}$, and thus, it can be approximated by the functions in $\mathcal{U}_{n}$ with respect to the pointwise convergence.
It remains to verify (iv) $\Rightarrow$ (v). To see this, suppose that $X\geq_{n}Y$. It follows from Lemma 1 that $\mathbb{E}[X]\geq\mathbb{E}[Y]$. First, we assume that $\mathbb{E}[X]>\mathbb{E}[Y]$, and the case that $\mathbb{E}[X]=\mathbb{E}[Y]$ will be studied later. Using Lemma 1 again and noting that $\mathbb{E}[X]>\mathbb{E}[Y]$, there exists $M\geq b$ such that $F_{X}^{[k]}(M)\leq F_{Y}^{[k]}(M)$ for all $k\in[n-1]$. For $u\in\mathcal{U}_{n}$, using integration by parts yields

### v4

@@ 1 Introduction
Stochastic dominance is a widely used concept in economics, finance, and decision-making under uncertainty, providing a robust method for comparing distributions of uncertain outcomes. This concept is essential in evaluating risk preferences without relying on a specific utility function or preference model, which allows for broad applications across various fields (Levy (2015); Peter (2021); Shaked and Shanthikumar (2007)).
@@
Despite its widespread use, the definition of higher-order stochastic dominance lacks consistency across the literature, sometimes leading to interpretational ambiguity. Consider, for example, two distributions, $F$ and $G$, each supported over the interval $[0,1]$. At first glance, one might assume that the question of whether $F$ dominates $G$ in fourth-order stochastic dominance would yield a straightforward answer. However, the consequence can depend significantly on the choice of the reference interval. For instance, if we assess the dominance using only the interval $[0,1]$, $F$ may not dominate $G$. Yet, extending the interval to $[0,2]$ might flip the assessment, resulting in $F$ dominating $G$ [^1] For a detailed discussion, see Example 1, where the specific distributions of $F=(8/9+\epsilon)\delta_{2/9}+(1/9-\epsilon)\delta_{1}$ and $G=\delta_{0}/3+2\delta_{4/9}/3$ with $\epsilon=1/100$. This highlights a crucial aspect of higher-order stochastic dominance: It can vary with alterations in the interval considered. This issue has led to ambiguous formulations of higher-order stochastic dominance across various texts. For example, Definition 7 in Baiardi et al. (2020) and the related definitions in Section 2.3 of Denuit and Eeckhoudt (2010) both adopted an arbitrary interval that encompasses the support of the distribution, but the definition actually depends on the choice of the interval.
@@
We say that the two formulations are consistent if the ranking between distributions $F$ and $G$ supported in $[a,b]$ remains the same when assessed under $n{\rm SD}_{\mathbb{R}}$ or $n{\rm SD}_{[a,b]}$. To the best of our knowledge, although various papers hint at the inconsistency issue under different settings (see the literature review below), the consistency of the two formulations of higher-order stochastic dominance was explicitly discussed only in Fang and Post (2022). In their Section 2.2, they contended that $n{\rm SD}_{[a,b]}$ imposes a more stringent criterion than $n{\rm SD}_{\mathbb{R}}$, suggesting that inconsistencies might arise when $n\geq 4$. We formally encapsulate these observations in our Proposition 1, providing a detailed analysis and illustrating the inconsistencies for cases where $n\geq 4$ with a straightforward counterexample in Example 1.
@@
One implication of our results is that, since stochastic domination relations (with $n\geq 4$) get strictly stronger when the reference interval shrinks, it affects both their applications and characterization results. For instance, a stochastic dominance relation is easier to hold when we enlarge the reference interval, which are usually harmless for real-data applications. The results also illustrate a drawback of the higher-order stochastic dominance relations. As stochastic dominance is mostly used as a robust tool for ordering risks without assuming specific preferences, the fact that they depend on a reference interval — a subjective choice of the modeler — jeopardizes their robustness interpretation. From a theoretical standpoint, it is sensible to introduce the reference interval under the assumption that it encompasses the support of all relevant random variables, allowing for a uniform analysis based on higher-order stochastic dominance with respect to this fixed reference interval. This is not the case when these theoretical results are applied to real-world problems such as portfolio selection and precautionary saving, where the support of the random variables often cannot be objectively specified, making the choice of the reference interval flexible.
Literature on stochastic dominance on grids and sub-intervals
Many studies have explored stochastic dominance for distributions restricted to specific subsets of $\mathbb{R}$, with several exploring the consistency of these orderings across various subsets. We discuss some of them here.
Fishburn (1976) investigated stochastic dominance on a restricted interval $[0,b]$, defined such that $X$ dominates $Y$ if $F_{X}^{[n]}(\eta)\leq F_{Y}^{[n]}(\eta)$ for all $\eta$ within $[0,b]$. [^2] The order $n$ can take any real value from $[1,\infty)$; however, for our purposes, we consider only integer $n$. Compared to $n{\rm SD}_{[0,b]}$, Fishburn’s criterion does not require the boundary condition at $b$, making it less stringent than both $n{\rm SD}_{[0,b]}$ and $n{\rm SD}_{\mathbb{R}}$. Furthermore, Fishburn (1980) showed that the stochastic dominance relations in Fishburn (1976) align with $n{\rm SD}_{\mathbb{R}}$ only for $n\in\{1,2\}$.333 For $n\in\{1,2\}$, the boundary condition that $F_{X}^{[n]}(b)\leq F_{Y}^{[n]}(b)$ guarantees that $F_{X}^{[n]}(\eta)\leq F_{Y}^{[n]}(\eta)$ for all $\eta>b$. In particular, for $\eta\geq b$, $F_{X}^{[1]}(\eta)=F_{Y}^{[1]}(\eta)=1$, and the comparison of $F_{X}^{[2]}(\eta)$ and $F_{Y}^{[2]}(\eta)$ reduces to comparing the expectations of $X$ and $Y$. However, this is not the case when $n\geq 3$ as higher moments appear. In light of our Proposition 1, we conclude that $n{\rm SD}_{[0,b]}$, $n{\rm SD}_{\mathbb{R}}$, and Fishburn’s criterion exhibit consistent for $n\in\{1,2\}$. However, for orders $n\geq 3$, these criteria do not exhibit consistency.
@@ 2 Main results
In this section, we will present all the results, while the proofs will be provided in the next section. For $a,b\in[-\infty,\infty]$ with $a<b$, denote by $\mathcal{X}_{[a,b]}$ the set of all bounded random variables taking values in $[a,b]$. For simplicity, we write $\mathcal{X}:=\mathcal{X}_{[-\infty,\infty]}$. We use capital letters, such as $X$ and $Y$, to represent random variables, and $F$ and $G$ for distribution functions. For $X\in\mathcal{X}$, we write $\mathbb{E}[X]$ for the expectation of $X$. Let $F_{X}$ denote the distribution function of $X$. We use $\delta_{\eta}$ to represent the point-mass at $\eta\in\mathbb{R}$. For a real-valued function $f$, let $f_{-}^{\prime}$ and $f^{\prime}_{+}$ be the left and right derivative of $f$, respectively, and denote by $f^{(n)}$ the $n$th derivative for $n\in\mathbb{N}$. Whenever we use the notation $f_{-}^{\prime}$, $f^{\prime}_{+}$ and $f^{(n)}$, it is understood that they exist. Denote by $[n]:=\{1,\dots,n\}$ with $n\in\mathbb{N}$. In this paper, all terms like “increasing”, “decreasing”, “convex”, and “concave” are in the weak sense.
@@
It is well-known that $F_{X}^{[n]}(\eta)$ is connected to the expectation of $(\eta-X)_{+}^{n-1}$ (see e.g., Proposition 1 of Ogryczak and Ruszczyński (2001)):
@@
[Shaked and Shanthikumar (2007, Section 4.A.7)] Let $n\in\mathbb{N}$. For $X,Y\in\mathcal{X}$, we say that $X$ dominates $Y$ in the sense of $n$th-order stochastic dominance on $\mathbb{R}$ ($n{\rm SD}_{\mathbb{R}}$), denoted by $X\geq_{n}Y$ or $F_{X}\geq_{n}F_{Y}$ if
@@
[Jean (1980, page 152)] Let $a,b\in[-\infty,\infty]$ with $a<b$ and $n\in\mathbb{N}$. For $X,Y\in\mathcal{X}_{[a,b]}$, we say that $X$ dominates $Y$ in the sense of $n$th-order stochastic dominance on $[a,b]$ ($n{\rm SD}_{[a,b]}$), denoted by $X\geq_{n}^{[a,b]}Y$ or $F_{X}\geq_{n}^{[a,b]}F_{Y}$ if
@@
$\mathbb{E}[(\eta-X)_{+}^{n-1}]\leq\mathbb{E}[(\eta-Y)_{+}^{n-1}],~{}\forall\eta\in[a,b]~{}{\rm and}~{}\mathbb{E}[(b-X)^{k-1}]\leq\mathbb{E}[(b-Y)^{k-1}]~{}{\rm for}~{}k\in[n].$
In contrast to $n{\rm SD}_{\mathbb{R}}$, $n{\rm SD}_{[a,b]}$ depends on a reference interval and involves additional boundary conditions that $F_{X}^{[k]}(b)\leq F_{Y}^{[k]}(b)$ of order $k\in[n]$. For $n\in[4]$, $n$SD corresponds to the well-known first, second, third and fourth-order stochastic dominance. Risk aversion, which includes aversion to mean-preserving spreads, aligns with second-order stochastic dominance as described in Rothschild and Stiglitz (1970). Higher orders of stochastic dominance, specifically third and fourth orders, cater to decision makers with more refined risk preferences. Third-order stochastic dominance reflects prudence (Kimball (1990)), while fourth-order dominance corresponds to temperance (Kimball (1992)). The characterizations of these preferences, as detailed by Eeckhoudt and Schlesinger (2006), extend from the traditional concept of mean-preserving spreads to a broader framework of risk apportionment.
For two random variables $X,Y\in\mathcal{X}_{[a,b]}$, both $n{\rm SD}_{\mathbb{R}}$ and $n{\rm SD}_{[a,b]}$ can be used to rank their order. This raises a natural question of whether these ranking relations are consistent. To examine this, we define two sets of utility functions that are regular $n$-increasing concave (Denuit and Eeckhoudt (2013)) over different domains:444 From the definition in Liu (2014), the utility functions in $\mathcal{U}_{n}$ exhibit $k$th-degree risk aversion for $k\in[n]$. Letting $n\to\infty$, $\mathcal{U}_{\infty}$ becomes the set of all completely monotone functions, which is well studied in the mathematics literature and is closely linked to Laplace–Stieltjes transforms (see, e.g., Schoenberg (1938)). In this case, utility functions express mixed risk aversion, as discussed in Caballé and Pomansky (1996).
@@
A direct explanation for why the implication from (iv) to (i) breaks when the order changes from 3 to 4 is provided below. Note that $n{\rm SD}_{[a,b]}$ involves additional boundary conditions compared to $n{\rm SD}_{\mathbb{R}}$. For $n=3$, suppose that $X\geq_{n}Y$. To check whether $X$ dominates $Y$ under $n{\rm SD}_{[a,b]}$, it is sufficient to verify only one boundary condition: $\mathbb{E}[(b-X)]\leq\mathbb{E}[(b-Y)]$, which is equivalent to $\mathbb{E}[X]\geq\mathbb{E}[Y]$. This comparison of expectations can be implied from $X\geq_{n}Y$. However, for $n=4$, the boundary condition of order $2$ arises: $\mathbb{E}[(b-X)^{2}]\leq\mathbb{E}[(b-Y)^{2}]$. In this case, $X\geq_{n}Y$ is not sufficient to establish the boundary condition. To illustrate this, we present an example involving two random variables $X,Y\in\mathcal{X}_{[0,1]}$ such that $X\geq_{4}Y$ and $\mathbb{E}[(1-X)^{2}]>\mathbb{E}[(1-Y)^{2}]$, serving as a counterexample to the implication (iv) $\Rightarrow$ (i) in Proposition 1 for $n=4$.
@@
Below we consider the rank of $X,Y$ under fourth-order stochastic dominance with different reference intervals. First, we consider $[0,1]$ as the reference interval. We have the conclusion that $X\not\geq_{4}^{[0,1]}Y$ as the following boundary condition with $k=3$ in Definition 2 is violated:
@@
Next, the reference interval is set as the entire real line, meaning that we consider fourth-order stochastic dominance as defined in Definition 1.
@@
By standard calculation, one can check that $\mathbb{E}[\left(\eta-X\right)_{+}^{3}]\leq\mathbb{E}[\left(\eta-Y\right)_{+}^{3}]$ for all $\eta\in\mathbb{R}$ (see Figure 1 for an intuitive illustration), which implies that $X\geq_{4}Y$. Finally, we set $[0,2]$ as the reference interval and aim to establish that $X\geq_{4}^{[0,2]}Y$. Note that $\mathbb{E}[\left(\eta-X\right)_{+}^{3}]\leq\mathbb{E}[\left(\eta-Y\right)_{+}^{3}]$ for all $\eta\in\mathbb{R}$. By Definition 2, it remains to verify the boundary condition with $k\in\{2,3\}$. By standard calculation, we have
$\mathbb{E}\left[\left(2-X\right)\right]-\mathbb{E}\left[\left(2-Y\right)\right]=-\frac{37}{8100}<0~{}~{}{\rm and}~{}~{}\mathbb{E}\left[\left(2-X\right)^{2}\right]-\mathbb{E}\left[\left(2-Y\right)^{2}\right]=-\frac{13}{2916}<0.$
Hence, the relation $X\geq_{4}^{[0,2]}Y$ holds.
@@ 3.1 Proof of Proposition 1
Note that $\overline{\mathcal{U}}_{n}^{[a,b]}$ is the closure of ${\mathcal{U}}_{n}^{[a,b]}$ with respect to the pointwise convergence. Thus, we only need to consider the statements (i), (ii), (iv) and (v). We aim to prove Proposition 1 by considering three steps in the following.
@@
Before showing the proof, we present an auxiliary lemma that will be used in all above steps.
@@
(iv) $\Rightarrow$ (i): Suppose that $X\geq_{3}Y$. It implies that $\mathbb{E}[(\eta-X)^{2}_{+}]\leq\mathbb{E}[(\eta-Y)^{2}_{+}]$ for all $\eta\in[a,b]$. By Lemma 1, we have $\mathbb{E}[X]\geq\mathbb{E}[Y]$, and thus, $\mathbb{E}[(b-X)]\leq\mathbb{E}[(b-Y)]$. Hence, we have concluded that $X\geq_{3}^{[a,b]}Y$, and this completes the proof.
@@
The implication (ii) $\Rightarrow$ (v) is trivial. The implication (v) $\Rightarrow$ (iv) is supported by the fact that the mapping $x\mapsto-(\eta-x)_{+}^{n-1}$ is contained in $\overline{\mathcal{U}}_{n}$, and thus, it can be approximated by a subset of functions in $\mathcal{U}_{n}$ with respect to the pointwise convergence. The implication (i) $\Leftrightarrow$ (ii) has been verified in Theorem 1 of Eeckhoudt et al. (2009). It remains to verify (iv) $\Rightarrow$ (v). In fact, this implication can be verified by the insight that every $u\in\mathcal{U}_{n}$ is a positive linear combination of singularity functions in the set $\{f:x\mapsto-(\eta-x)_{+}^{n-1}|\eta\in\mathbb{R}\}$; see Williamson (1956). We provide a self-contained proof based on the integration by parts below.
Suppose that $X\geq_{n}Y$. It follows from Lemma 1 that $\mathbb{E}[X]\geq\mathbb{E}[Y]$. First, we assume that $\mathbb{E}[X]>\mathbb{E}[Y]$, and the case that $\mathbb{E}[X]=\mathbb{E}[Y]$ will be studied later. Using Lemma 1 again and noting that $\mathbb{E}[X]>\mathbb{E}[Y]$, there exists $M\geq b$ such that $F_{X}^{[k]}(M)\leq F_{Y}^{[k]}(M)$ for all $k\in[n-1]$. For $u\in\mathcal{U}_{n}$, using integration by parts yields